Reports 1-1 of 1 Clear search Modify search
Detector Operation (Detector Operation)
zaza - 22:56 Friday 21 November 2025 (68215) Print this report
Operator Report - Afternoon shift

15:23 UTC  ITF in LN3 COMMISSIONING mode
15:40 UTC CALIBRATION MODE (Verkindt), unlock at 15:51 UTC. Ended at 17:37 UTC
17:37 UTC COMMISSIONING mode for ENV noise injections. Two unlocks at 17:59 UTC and 19:17 UTC. Tringali edited ENV_MAIN script to avoid the injections that had been successfully done before these two unlocks.
20:10, 20:44 UTC unlocks while waiting for the sensitivity curve to smoothen in order to try the injections a third time. B4_Cam1 showed irregularities, BS_IP, Seismon, Sea Activity red, troubles achieving LOCKED_ARMS_BEAT_DRMI_1F: injections activity stopped
21:31 UTC ITF left in LN3 SCIENCE mode with AUTOSCIENCE_ON, range oscillating between 40 and 52 Mpc
 

Images attached to this report
Comments to this report:
Tringali, Zaza - 0:14 Saturday 22 November 2025 (68219) Print this report

Comments on two unlocks during ENV injections:

  • the first unlock occurred during the magnetic injection performed in CEB (Figure 1),

  • the second unlock happened during the magnetic injection at WEB, after completing the injections at CEB and NEB (Figure 2).

It was not possible to complete the WEB magnetic injection nor to proceed with the planned acoustic injections.

The output files of CEB and NEB magnetic injections are collected in virgoData/NoiseInjections/MagneticInjectionsO4/output/:

MagneticSweep_NOISE_MAG_CEB-1447786253.txt
MagneticSweep_NOISE_MAG_NEB-1447787183.txt
 


 

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×